Suppose that the price of whiskey increases from R100 to R150 a bottle and as a result the quantity demanded decreases from 1 100 bottles to 800 bottles
.Use the income elasticity of demand to distinguish between a normal good and an inferior good. In your explanation, provide the correct elasticity coefficient for each product and the relationship between income and quantity demanded. (4)
5.3. The cross-price elasticity coefficient between Apple and Android is 2,0. Are these goods complements or substitutes
1
Expert's answer
2021-04-19T07:36:16-0400

Solution:

Elasticity of demand = "\frac{\%\;change\; in\; quantity\; demanded}{\%\; change\; in\; price}"


% change in qty demanded = "\frac{Q_{2} -Q_{1}}{(Q_{2}+Q_{1})/2 } \times 100 = \frac{800 -1100}{(800+1100)/2 } \times 100"


"\frac{-300}{950} \times 100 = -0.316\times 100 = -31.6\%"


% change in price ="\frac{P_{2} -P_{1}}{(P_{2}+P_{1})/2 } \times 100 = \frac{150 -100}{(150+100)/2 } \times 100"


"\frac{50}{125} \times 100 = 0.4\times 100 = 40\%"


Elasticity of demand = "\frac{-31.6\%}{40\%} = 0.79"


The elasticity of demand is less than 1, which means that whiskey is price inelastic and a normal good.

 

A normal good is one whose demand rises as consumer income increases. On the other hand, an inferior good is a good whose demand decreases when consumer income rises, or increases when consumer income decreases.

The elasticity coefficient for a normal good: YED>0 (Income elasticity of demand is positive but less than one).

The elasticity coefficient for an inferior good: YED<0 (Income elasticity of demand is negative and less than one)

 

A cross elasticity of 2.0 means that Apple and Android are perfect substitutes. Since they have a high positive cross elasticity of demand.